# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

### v1

@@ B.4 Learning the Scoring Function
Because 4.2 and 4.1 hold with no error, by Lemma B.2 and Lemma B.1, it is easy to show that $\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)=\Pr(x_{i},x_{j}\mid z)\;\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$ and $\Pr(x_{i},z_{i}\mid z)=\Pr(x_{i}\mid z)\;\Pr(z_{i}\mid z)$.
@@
Appendix C A Heuristic Generalization to Complex Signals
@@
Suppose two agents review the same paper and each has a review after reading the paper, which is denoted as $X_{i}$ and $X_{j}$ respectively. Instead of working hard to obtain a high-quality review, each agent can use an LLM to generate a fake review, denoted as $Z_{i}$ and $Z_{j}$. The first challenge is how to measure the level of agreement between two reviews. The idea is to represent each review with an embedding vector and compute the similarity between a pair of embedding vectors. Depending on the information of interest, different embedding methods can be applied. For example, classic NLP embedding methods such as Word2Vec [Mikolov, 2013], TF-IDF [Manning, 2009], BERT and its variants [Devlin, 2018, Sanh, 2019], and GPT-3 [Brown et al., 2020] can be used to represent the language-level information within the responses.
@@
The same idea can be generalized beyond peer review. Suppose the principal has a pre-defined set of questions about the responses, denoted as $\mathcal{Q}$ where $|\mathcal{Q}|=K$. Each question requires a (normalized) score between $-1$ and $1$, ensuring that the inner product between vectors can serve as a meaningful similarity measure. Given a response $X_{i}$, the principal can prompt an LLM to iteratively get an answer to each question $q\in\mathcal{Q}$. This process converts each response $X_{i}$ into a vector $\phi(X_{i})$ of length $K$, and each entry is a score between $-1$ and $1$. Similar ideas have been applied to identify similar data points [Zeng et al., 2024], analyze peer review [Liang et al., 2024], and select features [Jeong et al., 2024].
@@
By projecting the embedding vectors onto the orthogonal complement of $\phi(Z)$, we remove the information that already exists in $Z$. This projection approach is also used in NLP for “debiasing” word embeddings by eliminating similarity along a “biased direction” [Bolukbasi et al., 2016].
@@
Appendix D Experiments (Continue)
D.1 Assumption Verifications (Continue)
Figure 5 present the TVD mutual information between LLM samples and human responses conditioned on (different) LLM samples, $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$, which is the analogous results to Fig. 2 on the remaining datasets.
(a) Hatefulness
(b) Offensiveness
Figure 5: Analogous figures to Fig. 2 for other datasets.
Table 2: TVD-Conditioned Mutual Information Between Two LLMs on the Preference Alignment Data
@@ D.1 Assumption Verifications (Continue)
Table 2, Table 3, Table 4 and Table 5 present the mutual information $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ on four datasets. $Z$ and $Z_{i}$ are sampled by the same LLM indicated on the rows and $Z_{j}$ is sampled by the LLM indicated on the columns. The first observation is that all values displayed in black exceed $0.05$, sometimes significantly so, which represents the mutual information of the random reporting baseline. This implies that $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)>\epsilon$ holds consistently, rejecting hypothesis 2.
Regarding hypothesis 4, we observe that the average correlations between two LLMs are generally weaker than those between two human labelers, provided $Z_{i}$ and $Z$ are independently generated by the same model. This trend is evident in the tables, where almost all bracketed values are positive, with one exception.***The exception occurs in the preference alignment dataset, where samples of $Z_{i},Z_{j}$, and $Z$ all generated by GPT-4.
Table 3: TVD-Conditioned Mutual Information Between Two LLMs on the Hatefulness Labeling Data
@@
Table 4: TVD-Conditioned Mutual Information Between Two LLMs on the Offensiveness Labeling Data
@@
Table 5: TVD-Conditioned Mutual Information Between Two LLMs on the Toxicity Labeling Data
@@
D.2 Detecting Cheating Agents (Continue)
D.2.1 Additional Figures and Tables
In Fig. 6, we present additional figures on the toxicity dataset, analogous to Fig. 4(a), for other types of LLM cheaters. Analogous results to Table 1 for other datasets are summarized in Table 6, Table 7, Table 8, and Table 9. These results support our main claim that, although our method is occasionally slightly outperformed by certain baselines, it consistently demonstrates the most robust performance across all tested settings.
@@ D.2.1 Additional Figures and Tables
Figure 6: The area under the ROC curve (AUC) for various methods while varying the fraction of LLM cheaters. These are the analogous figures for Fig. 4(a) for the remaining LLMs on the toxicity labeling dataset.
Table 6: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-3.5 cheaters.
@@
Table 7: Average AUC (and bottom 10% quantiles) across datasets and methods with Mistral cheaters.
@@
Table 8: Average AUC (and bottom 10% quantiles) across datasets and methods with Gemma cheaters.
@@
Table 9: Average AUC (and bottom 10% quantiles) across datasets and methods with Phi cheaters.
@@
D.2.2 Detection Performance vs. Number of Samples (Continue)
Figure 7 presents similar results to Fig. 4(b) for the hatefulness, offensiveness, and toxicity labeling datasets.
(a) Hatefulness
(b) Offensiveness
(c) Toxicity
@@
D.2.3 Conditioning on samples of multiple LLMs
In practice, agents may employ a variety of LLMs for low-effort response generation, and the principal lacks prior knowledge of which models they are using. To address this challenge, we propose an iterative conditioning approach, systematically removing the influence of responses generated by each LLM available to the principal. The idea is to iteratively compute the CA score conditioned on each of the LLM samples the principal has, and then take the minimum as the final score. Intuitively, human agents will perform consistently well when conditioning out any of the LLMs’ information. However, LLM cheaters will perform poorly in at least one of the iterations and thus lead to a lower final score.
In Fig. 8, we replicate Fig. 4(a) for the setting where the cheaters randomly pick two LLMs to cheat with half-half probability while the principal iteratively conditions out his samples for these two LLMs’ answers. Our method exhibits the most robust performance across all settings, echoing the previous conclusion.
We further note that the problem of LLM-assisted cheating is less of a concern when each cheater uses a unique LLM. This is because LLM responses are less correlated with each other and can be easily detected as outliers under classic crowdsourcing evaluation methods, such as the EM algorithm under the DS model.
(a) Mistral & GPT-3.5
(b) Mistral & GPT-4
@@ D.2.3 Conditioning on samples of multiple LLMs
Figure 8: Area under the ROC curve (AUC) for various methods as the fraction of LLM cheaters increases. In each case, LLM cheaters randomly choose between two LLMs to generate responses, and the score is defined as the minimum of the two conditioned CA scores, each conditioned on samples from one of the two LLMs.
D.3 Complex Signals
We further test the idea illustrated in Appendix C that generalizes our method to high-dimensional settings.
In this experiment, we randomly sample 500 papers from the ICLR 2024 OpenReview data and select three human reviews per paper as human responses $X$. For each of the same batch of papers, we further use each of the five LLMs to generate three AI reviews using three similar-meaning yet different prompts (Appendix E). We randomly pick one of the generated reviews as the principal’s sample $Z$, and the other two generated reviews will be used to simulate LLM-cheaters’ responses.
We prompt GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector with the prompt shown in Appendix E. We further simulate a review embedding tensor with a size of $n\times m\times 19$, where there are $n=188$ reviewers and $m=500$ papers in total. Each reviewer randomly reviews $k\in\{5,8\}$ papers. An $\alpha_{llm}$ fraction of reviewers are LLM cheaters whose review vectors are replaced with a particular type of LLM-generated reviews. An $\alpha_{p}$ fraction of reviewers are prior cheaters, where the score in each entry of the review vector is randomly sampled according to the marginal distribution in the dataset. An $\alpha_{r}$ fraction of reviewers are random cheaters, where the score in each entry of the review vector is sampled uniformly at random. In our experiments, we independently sample each of $\alpha_{llm}$, $\alpha_{p}$, and $\alpha_{r}$ uniformly from the interval $[0,0.1]$ and take average over $100$ trials. While constructing the review embedding tensor, we guarantee that each paper has approximate $3$ reviews so that the variance of the number of reviewers per paper is controlled. Again, we consider five types of LLMs and for each type of LLM cheater, the principal’s sample $Z$ is an independent generation of the same LLM under a different prompt.
@@ D.3 Complex Signals
  • Projected Cosine Similarity: Applies the method described in Appendix C, where each review vector is projected onto the subspace orthogonal to the principal reviewer’s sampled embedding before computing similarity.
@@
We present the results in Fig. 9. It is clear that conditioning out the principal’s sampled LLM reviews is more effective than computing the cosine similarity directly. Furthermore, naively removing reviews that are similar to $Z$ has a large variance, whose performance greatly depends on the existence of random and prior cheaters.
@@
Appendix E Prompts Templates
The prompt templates for two different dataset settings are provided in Table E and Table E, respectively.

### v2

@@ Appendix E Learning the Scoring Function
Because 4.2 and 4.1 hold with no error, by Lemma D.2 and Lemma D.1, it is easy to show that $\Pr(x_{i},x_{j},z_{i},z_{j}\mid z)=\Pr(x_{i},x_{j}\mid z)\;\Pr(z_{i}\mid z)\;\Pr(z_{j}\mid z)$ and $\Pr(x_{i},z_{i}\mid z)=\Pr(x_{i}\mid z)\;\Pr(z_{i}\mid z)$.
@@
Appendix F A Heuristic Generalization to Complex Signals
@@
Suppose two agents review the same paper and each has a review after reading the paper, which is denoted as $X_{i}$ and $X_{j}$ respectively. Instead of working hard to obtain a high-quality review, each agent can use an LLM to generate a fake review, denoted as $Z_{i}$ and $Z_{j}$. The first challenge is how to measure the level of agreement between two reviews. The idea is to represent each review with an embedding vector and compute the similarity between a pair of embedding vectors. Depending on the information of interest, different embedding methods can be applied. For example, classic NLP embedding methods such as Word2Vec mikolov2013efficient, TF-IDF manning2009introduction, BERT and its variants devlin2018bert, sanh2019distilbert, and GPT-3 brown2020language can be used to represent the language-level information within the responses.
@@
The same idea can be generalized beyond peer review. Suppose the principal has a pre-defined set of questions about the responses, denoted as $\mathcal{Q}$ where $|\mathcal{Q}|=K$. Each question requires a (normalized) score between 1-1 and $1$, ensuring that the inner product between vectors can serve as a meaningful similarity measure. Given a response $X_{i}$, the principal can prompt an LLM to iteratively get an answer to each question $q\in\mathcal{Q}$. This process converts each response $X_{i}$ into a vector $\phi(X_{i})$ of length $K$, and each entry is a score between 1-1 and $1$. Similar ideas have been applied to identify similar data points zeng2024similar, analyze peer review liang2024mapping, and select features jeong2024llmselect.
@@
By projecting the embedding vectors onto the orthogonal complement of $\phi(Z)$, we remove the information that already exists in $Z$. This projection approach is also used in NLP for “debiasing” word embeddings by eliminating similarity along a “biased direction” bolukbasi2016man.
@@
Appendix G Experiments (Continue)
G.1 Large Language Models
In this work, we select five well-known LLMs to generate labels or annotations for all used datasets, including GPT-3.5-turbo, GPT-4, Gemma-2-2b-it [^3] https://huggingface.co/google/gemma-2-2b-it, Phi-3.5-mini-Instruct  [^4] https://huggingface.co/microsoft/Phi-3.5-mini-instruct, and Mistral-7B-Instruct-v0.3555https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3. The corresponding prompt templates are provided in Appendix H. Note that we choose only five models because other commonly used LLMs such as LLaMA and Gemini refuse to answer toxicity-related questions.
G.2 Details of the Baseline Methods
Output Agreement (OA)  Perhaps the most straightforward idea to score an agent is to count the frequency of agreements between her responses and another agents’ responses. Specifically, while scoring agent $i$, we iteratively pair her with a different agent $j$ and let $M_{i,j}$ denote the set of questions answered by both agents. Then, the score of agent $i$ is the average of the empirical probability of agreement, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}[\hat{X}_{i,q}=\hat{X}_{j,q}]}{|M_{i,j}|}$ where $\hat{X}$ is the crowdsourced dataset with rows representing agents and columns representing questions. [^6] Let $\frac{0}{0}=0$. A score of 1 thus indicates perfect agreement while a score of 0 indicates no agreement at all.
Correlated Agreement (CA)  We use the CA mechanism shnayder2016informed as a baseline which does not incorporate the principal’s information $Z$. Details of the mechanism have been presented in Algorithm 1.
Expectation-Maximization (EM)  The EM algorithm is a well-known approach to estimating the reliabilities of noisy agents and better recovering the ground truth considering the reliabilities of agents [dawid1979maximum]. Each agent $i$ is assumed to make mistakes according to a confusion matrix $\Gamma^{i}$ where $\Gamma^{i}_{h,l}$ denotes the probability that agent $i$ reports $h$ when the ground truth answer is $l$. The EM algorithm first initializes a prior of the ground truth for each question and a confusion matrix for each agent. Then, it iteratively goes through an E-step and an M-step. In the E-step, the algorithm computes the posterior of the ground truth of each question given the prior and the confusion matrix. In the M-step, the confusion matrix of each agent is updated using the posterior computed in the E-step. The above two steps are iterated until the likelihood of $D$ converges. Finally, we score the reliability of agent $i$ as $\sum_{h\in\Sigma}\Gamma^{i}_{h,h}\tilde{P}(h)$ where $\tilde{P}(h)$ is the empirical marginal distribution of observing $h$ in the dataset.
Conditioned Agreement ($\text{OA}_{Z}$)  The first three measures do not utilize the principal’s information about the cheap signal, denoted as $Z$. Note that in our experiments, $Z$ is the vector of labels generated by a particular LLM. The conditioned agreement metric scores a agent if she can provide agreements to another agent in addition to $Z$. Similar to the agreement metric, the score of agent $i$ is the average of the empirical probability of agreement conditioned on $Z$, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathds{1}[\hat{X}_{i,q}=\hat{X}_{j,q}\&\hat{X}_{i,q}\neq Z_{q}]}{|M_{i,j}|}$. Intuitively, if agent $i$ uses an LLM whose generated labels are similar to $Z$, then $\hat{X}_{i,q}\neq Z_{q}$ will happen with a small probability, meaning that the conditioned agreement score will be small.
G.3 Assumption Verifications (Continue)
Table 2, Table 3, Table 4 and Table 5 present the mutual information $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ on four datasets. $Z$ and $Z_{i}$ are sampled by the same LLM indicated on the rows and $Z_{j}$ is sampled by the LLM indicated on the columns. The first observation is that all values displayed in black exceed $0.05$, which represents the mutual information of the random reporting baseline. This implies that $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)>\epsilon$ holds consistently, rejecting hypothesis 2.
Regarding hypothesis 4, we observe that the average correlations between two LLMs are generally weaker than those between two human agents, provided $Z_{i}$ and $Z$ are independently generated by the same model. This trend is evident in the tables, where almost all bracketed values are positive, with one exception. [^7] The exception occurs in the preference alignment dataset, where samples of $Z_{i},Z_{j}$, and $Z$ all generated by GPT-4.
Table 2: Conditional TVD Mutual Information Between Two LLMs on the Preference Alignment Data
@@ G.3 Assumption Verifications (Continue)
Table 3: Conditional TVD Mutual Information Between Two LLMs on the Hatefulness Labeling Data
@@
Table 4: Conditional TVD Mutual Information Between Two LLMs on the Offensiveness Labeling Data
@@
Table 5: Conditional TVD Mutual Information Between Two LLMs on the Toxicity Labeling Data
@@
G.4 Detecting Low-effort Agents (Continue)
G.4.1 Additional Figures and Tables
@@
Figure 4: The area under the ROC curve (AUC) for various methods while varying the fraction of LLM-reliant agents. These are the analogous figures for Figure 3(b) for the remaining LLMs on the toxicity labeling dataset.
Table 6: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-3.5-reliant agents.
@@
Table 7: Average AUC (and bottom 10% quantiles) across datasets and methods with Mistral-reliant agents.
@@
Table 8: Average AUC (and bottom 10% quantiles) across datasets and methods with Gemma-reliant agents.
@@
Table 9: Average AUC (and bottom 10% quantiles) across datasets and methods with Phi-reliant agents.
@@
In Figure 4, we present additional figures on the toxicity dataset, analogous to Figure 3(b), for other LLMs. Analogous results to Table 1 for other datasets are summarized in Table 6, Table 7, Table 8, and Table 9. These results support our main claim that, although our method is occasionally slightly outperformed by certain baselines, it consistently demonstrates the most robust performance across all tested settings.
G.4.2 Detection Performance vs. Number of Samples
We further examine how many questions the principal needs to sample to effectively detect low-effort agents. In Figure 5, we report the AUC of the conditioned CA mechanism with the fraction of LLM-reliant agents fixed at $0.15$, while the fractions of random and biased agents are independently sampled from $[0,0.2]$. We vary the sampling rate from 20% to 100%. The results vary across the datasets and the LLMs that agents use. Except for the hatefulness labeling dataset with GPT-3.5-reliant agents, sampling 50% of the questions is sufficient to outperform the baseline CA mechanism.
(a) Preference
(b) Hatefulness
(c) Offensiveness
(d) Toxicity
@@
G.4.3 Conditioning on samples of multiple LLMs
In practice, agents may employ a variety of LLMs for low-effort response generation, and the principal lacks prior knowledge of which models they are using. To address this challenge, we propose an iterative conditioning approach, systematically removing the influence of responses generated by each LLM available to the principal. The idea is to iteratively compute the CA score conditioned on each of the LLM samples the principal has, and then take the minimum as the final score. Intuitively, human agents will perform consistently well when conditioning out any of the LLMs’ information. However, LLM-reliant agents will perform poorly in at least one of the iterations and thus leading to a lower final score.
In Figure 6, we replicate Figure 3(b) for the setting where the LLM-reliant agents randomly pick two LLMs with half-half probability while the principal iteratively conditions out his samples for these two LLMs’ answers. Our method exhibits the most robust performance across all datasets, echoing the previous conclusion.
(a) Gemma & GPT-3.5
(b) Gemma & GPT-4
@@ G.4.3 Conditioning on samples of multiple LLMs
Figure 6: Area under the ROC curve (AUC) for various methods as the fraction of LLM-reliant agents increases. In each case, LLM-reliant agents randomly choose between two LLMs to generate responses, and the score is defined as the minimum of the two conditioned CA scores, each conditioned on samples from one LLM.
G.5 Complex Signals
We further test the idea illustrated in Appendix F that generalizes our method to high-dimensional settings.
In this experiment, we randomly sample 500 papers from the ICLR 2024 OpenReview data and select three human reviews per paper as human responses $X$. For each of the same batch of papers, we further use each of the five LLMs to generate three AI reviews using three similar-meaning yet different prompts (Appendix H). We randomly pick one of the generated reviews as the principal’s sample $Z$, and the other two generated reviews will be used to simulate the responses of LLM-reliant agents.
We prompt GPT-4 to process each human review and LLM-generated review into a $19$-dimension vectors with the prompt shown in Appendix H. We further simulate a review embedding tensor with a size of $n\times m\times 19$, where there are $n=188$ reviewers and $m=500$ papers in total. Each reviewer randomly reviews $k\in\{5,8\}$ papers. An $\alpha_{llm}$ fraction of reviewers are LLM-reliant agents whose review vectors are replaced with a particular type of LLM-generated reviews. An $\alpha_{p}$ fraction of reviewers are prior-reporting agents, where the score in each entry of the review vector is randomly sampled according to the marginal distribution in the dataset. An $\alpha_{r}$ fraction of reviewers are random agents, where the score in each entry of the review vector is sampled uniformly at random. In our experiments, we independently sample each of $\alpha_{llm}$, $\alpha_{p}$, and $\alpha_{r}$ uniformly from the interval $[0,0.1]$ and take average over $100$ trials. While constructing the review embedding tensor, we guarantee that each paper has approximate $3$ reviews so that the variance of the number of reviewers per paper is controlled. Again, we consider five types of LLMs and for each type of LLM-reliant agent, the principal’s sample $Z$ is an independent generation of the same LLM under a different prompt.
@@ G.5 Complex Signals
  • Projected Cosine Similarity: Applies the method described in Appendix F, where each review vector is projected onto the subspace orthogonal to the principal reviewer’s sampled embedding before computing similarity.
@@
We present the results in Figure 7. It is clear that conditioning out the principal’s sampled LLM reviews is more effective than computing the cosine similarity directly. Furthermore, naively removing reviews that are similar to $Z$ has a large variance, whose performance greatly depends on the existence of random agents and prior-reporting agents.
@@
Appendix H Prompts Templates
We adopt our prompt templates based on recent studies [pang2024improving, liu2024automatic, pang2025token]. The specific templates used for the two dataset settings are presented in Table H and Table H, respectively.